OpenStudy (anonymous):

How can you represent the terms of a sequence explicitly? How can you represent them recursively? Provide mathematical examples to support your opinions.

OpenStudy (jamesj):

For example 1, 2, 3, 4, 5, ... An explicit way to write this would be \( a_n = n \). Now, what's a recursive definition for the same sequence?

OpenStudy (perl):

an+1= an + 1

OpenStudy (jamesj):

Yes, \[ a_{n+1} = a_n + 1 \] Now, you make up another example for yourself.

OpenStudy (perl):

a(n+1) = a(n) + 1 , a(1) = 1

OpenStudy (perl):

james, can you help me with work problem?

OpenStudy (jamesj):

actually yes, what perl wrote just now is right. It's very important to give the first case

OpenStudy (perl):

can you do integrals?

OpenStudy (jamesj):

yes, give me the link and I'll help if I can

OpenStudy (anonymous):

@JamesJ Question... What is the 15th term in the sequence using the given formula? cn = 3n – 1 Is 44 the right answer?

OpenStudy (anonymous):

Just wondering if I understood this all correctly.

OpenStudy (jamesj):

clearly, yes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Sorry, last question.. Is my recursive formula for 2,4,6,8,10,12,14,16,18,20. . . correct? a(n)=a(n-1)+2

OpenStudy (jamesj):

Looks right to me!